A wave has a frequency of 25 Hz. It has a wavelength of 60 m. What is the period and speed of the wave?
1
Expert's answer
2011-04-13T11:36:08-0400
You can use the following formula:
T = 1/f = 1/25 s.
C = λ/T = 60*25 = 1500 m/s